# Instruct-Tuning Pretrained Causal Language Models for Ancient Greek Papyrology and Epigraphy

### v1

@@ 4 Discussion
Table 1: Comparison of models for restoration, geographical attribution, and dating tasks. The results for Epigr_1_Llama3.1-8b-Inscruct are from this study, while results from Assael, Sommerschield et al. (2022) are included for comparison.
@@
Method CER (%) Top-1 (%) Top-20 (%) Top-1 (%) Top-3 (%) Years
Epigr_1_Llama3.1-8b-Inscruct 22.5 60.9 77.5 75.0 83.7 26.2
Assael et al. (2022) 26.3 61.8 78.3 70.8 82.1 29.3
Papy_1_Llama3.1-8b-Inscruct 16.3 71.3 85.0 66.5 80.0 21.7
Pavlopoulos et al. (2023) 54
Table 1 summarizes the results and allows for a comparison between the models developed in this study and the figures reported by Assael and colleagues [4]. Notably, the datasets used for testing Epigr_1_Llama3.1-8b-Instruct were smaller than those employed by Assael and colleagues [4], with 1,856 samples for dating and 2,870 for geographical attribution, compared to their dataset of 7,811 samples (though it is possible that not all of their entries in the test set included chronological or geographical data). A full, comparable replication of their study using a model like LLaMA would provide a more definitive assessment of which approach yields the best results. However, the primary aim of this experiment was not to maximize numerical performance. The goal was to determine whether a simple, instruct-based approach is competitive and warrants further exploration. The findings suggest that this is the case, and that additional experimentation with smaller, larger and more advanced similar base models, along with various approaches to dataset formatting, masking, training, and model-merging, would be worthwhile.
@@
The key takeaway is that the tools developed here at least approximate the performance of the state-of-the-art model. Additionally, they offer a significant advantage by being applicable when scholars have an approximate sense of the number of missing characters but lack precise information on the number of missing word boundaries—an issue commonly encountered in the study of ancient textual artifacts. Furthermore, the papyri models introduced in this study are the first specifically designed for handling such material.
One legitimate concern that arises is whether some of the texts in the test datasets might have been included in the trillions of tokens used during the pretraining of LLaMA 3.1 8B. Although these texts were not part of the fine-tuning dataset, it is possible that the model had previously encountered this data. To address this concern, the models were tested on two unedited papyri from Uppsala University Library, with 100 differently masked versions of each text generated for ad-hoc evaluation.
  • Pap.Ups 106 (see image at urn.kb.se/resolve?urn=urn:nbn:se:alvin:portal:record-188165
    and text in Appendix 1) was processed with a character error rate (CER) of 17.2%, a top-1 accuracy of 62.0%, and a top-20 accuracy of 84.0%. The model plausibly assigns the document to Oxyrhynchus and accurately dates it to 71 AD (the fourth year of Vespasian, 71–72 AD).
  • Pap.Ups 18 (see image at urn.kb.se/resolve?urn=urn:nbn:se:alvin:portal:record-187926
    and text in Appendix 1) is more heavily damaged, resulting in lower model performance, but still provided useful results with a CER of 28.9%, a top-1 accuracy of 55.0%, and a top-20 accuracy of 71.0%. The model plausibly places it in Oxyrhynchus and dates the fragment to 304 AD, approximately 11–12 years from the correct date of around 292 AD (the eighth year of Diocletian and the seventh of Maximian are mentioned).
For the reconstruction tasks, however, there were only insignificant deteriorations of 0.1–0.2% in CER and top-1 accuracy, but a 3% increase in top-20 accuracy. Hence, the model’s overall restoration capacity interestingly improved when it was re-infused with the base model.
@@ 5 Conclusion
This experiment demonstrated the effectiveness of fine-tuning pretrained causal language models for dating, geographic attribution, and text restoration in ancient Greek inscriptions and documentary papyri using a simple chat template. While the results are preliminary, they suggest that this simple approach not only comes close to but, in some respects, surpasses the current state of the art. The simplicity and scalability of this method, particularly the ease of fine-tuning newer models on the same data without much modification, is a key advantage. There is significant potential for further improvement through more refined data cleaning, data augmentation and masking strategies, optimized training parameters, and experimentation with smaller or larger and more advanced base models as they become available in the future.
Although more specialized models may eventually outperform this simple method, instruct-tuned pretrained causal language models can offer an efficient and easily implementable moving baseline. Researchers developing machine learning methods for textual criticism could either aim to surpass this baseline or incorporate it as part of a more comprehensive solution. Rather than viewing different models as being in competition, a more productive approach may involve combining them. A system that integrates several models, supported by an adjudicator or ranker model, could potentially achieve the best results. This collaborative approach, taking advantage of the complementary strengths of various forms of AI, holds great promise for developing powerful tools to advance philological research.

### v2

@@ 4 Discussion
Table 1: Results for restoration, geographical attribution, and dating.
@@
CER (%) Top-1 (%) Top-20 (%) Top-1 (%) Top-3 (%) Years
Epigr_2_Llama3.1-8b-Instruct on my test 20.5 63.7 83.0 75.0 83.7 37.1
Assael et al. (2022) on their test 26.3 61.8 78.3 70.8 82.1 29.3
Epigr_1_Llama3.1-8b-Instruct on shared test 22.6 61.4 81.6 65.0 78.4 55.7
Assael et al. (2022) on shared test set 26.3 61.8 78.3 69.0 80.2 48.3
Epigr_1_Llama3.1-8b-Instruct on new inscr. 30.8 48.9 72.9 - - -
Assael et al. (2022) on new inscr. 31.1 46.5 68.3 - - -
Papy_1_Llama3.1-8b-Instruct on my test 14.9 73.5 85.9 66.5 80.0 21.7
Papy_1_Llama3.1-8b-Instruct on new papyri 29.67 73.0 89.0 - - -
Pavlopoulos et al. (2023) on their test - - - - - 54
Table 1 summarizes the results, providing a preliminary comparison between the models developed in this study and Ithaca [4]. While a new experiment fine-tuning LLaMA using an 80/10/10 split would be necessary for a more accurate comparison, the superior results on the test set with new inscriptions suggest that the models developed here generalize well to unseen data, rather than merely overfitting to the test set. Furthermore, the results on unedited or recently edited papyri and inscriptions imply that the strength of these models does not simply stem from possible exposure to test texts included in the data used during Meta’s pretraining process, but rather from the effects of the fine-tuning process itself. Although they exhibit a higher CER and lower top-1 and top-20 accuracy on this material, the fact that Ithaca’s performance was similarly affected, along with the variation in results between the two papyri, suggests that the challenge lies in the task itself rather than a lack of prior exposure.
In any case, the primary goal of this experiment was to assess whether a simple, instruction-based approach could compete with more established methods in this area, and the findings indicate that it can. These results justify further exploration through additional experiments involving smaller, larger, and more advanced base models, as well as alternative strategies for dataset formatting, masking, training, and model merging. The tools developed in this study perform at least comparably to the state-of-the-art in text restoration, demonstrating significant potential. However, the method’s effectiveness for dating and geographic attribution was less conclusive.
@@
Moreover, these models offer a notable advantage: they are particularly useful in cases where scholars have an approximate sense of the number of missing characters but lack precise information about missing word boundaries—a frequent challenge in the study of ancient textual artifacts. Additionally, the papyri models introduced here represent a new resource for this material.
@@ 5 Conclusion
This experiment demonstrated the efficacy of fine-tuning pretrained causal language models for text restoration in ancient Greek inscriptions and documentary papyri using a simple instruction template. The method’s success in dating and geographic attribution was less definitive. While the results are preliminary, they suggest that this straightforward approach not only approaches but, in some respects, may surpass the current state of the art. The simplicity and scalability of this method, particularly the ease of fine-tuning newer models on the same data with minimal modification, is a key advantage. There is significant potential for further improvement through refined data cleaning, data augmentation, and masking strategies, optimized training parameters, and experimentation with smaller, larger, and more advanced base models as they become available.
Although more specialized models may eventually outperform this straightforward method, instruction-tuned pretrained causal language models can serve as an efficient and easily implementable moving baseline. Researchers developing machine learning methods for textual criticism could aim either to surpass this baseline or to incorporate it as part of a more comprehensive solution. Rather than viewing different models as competitors, a more productive approach might involve combining them as nodes in a system supported by an adjudicator or ranker model. This collaborative approach, drawing on the complementary strengths of various AI models, likely holds the greatest promise for developing powerful tools to support philological research.